# Graph Residual based Method for Molecular Property Prediction

### v1

@@ 3.4 End to End Learning
Since there are two problems to be solved, regression and classification, both will have their own independent learning schemes. The entire GNN backbone can be represented as $F_{n}:G(n)\rightarrow m_{f}$, where $F_{n}$ is the GNN function that generates the molecular fingerprint/embeddings $(m_{f})$. The MLP as defined in Sec.3.2 is automatically integrated in the GNN function which allows for backpropagation. The embeddings are then used for the general downstream tasks. For the regression task, we can define the task in the format $MLP(m_{f})\rightarrow\text{output prediction}$.
@@
The multilayer-perceptron is applied to molecular embeddings generated by the GNN. Since the entire updation takes place in a single setting and is bound by a single computation graph, proper end-to-end learning takes place. For the multi-class multi-label classification setting, a diagrammatic representation of the workflow has been represented in Figure 10. As can be seen, the SMILES data is first converted to a graph database by PyTorch-geometric [11], with the features of the nodes and edges being extracted by Rdkit [55]. The dataset is then passed through the GNN as represented above to generate $m_{f}$.The VAE part of the execution can be represented as $VAE(m_{f})\rightarrow\hat{p}$ and $\sigma(\hat{p})\rightarrow\text{predicted vector}$.
In the current scenario, an MLP is used as the encoder and the decoder of the VAE, however, it can be easily replaced with much more sophisticated VAE methods such as [56]. The VAE generates a probability vector which when passed through the sigmoid function will give us the predicted vector. In both the regression and classification scenarios, all the entities are attached to an individual computation graph. Thus, this allows for an end-to-end manner for training where the molecular fingerprint generated is dependent on all the entities and backpropagation is allowed throughout the computation graph.
@@ 4 Model Setup and Training
For the model setup, we use the GNN backbone described in Sec. 3.1 for both the regression and classification examples. All the hyper-parameters of the graph convolutions as well as the fully connected layers remain the same. In the case of regression property prediction, Huber loss has been used, a combination of MSE and MAE, to account for the presence of outliers. The resultant loss generated is converted into an exponential loss function.
In the case of Multi-class/Multi-label Classification datasets, the Cross entropy loss function has been used. The GNN backbone along with a sigmoid layer of one-hot encoded classes has been used. To prevent class imbalance from affecting our results, a self-adaptive weighted loss-function has been implemented controlled by Adam optimizer [57] is implemented. This prevents some classes from becoming over-dominant thus increasing our F1 score.
The GNN backbone, in conjunction with VAE, has been used to establish an end-to-end pipeline for training. The hyperparameters of the GNN backbone remain the same, and those of the VAE have been described in Sec. 3.3. To explain, the efficacy of the GNN-VAE model, we also present comparative results of the GNN backbone with the last layer as a sigmoid activation function.
Adam optimizer as well as Cosine annealing scheduler with warm restarts has been used in both the regression as well as classification cases. The batch size is 64, as it is found to be the optimum one, and we have run the model for a maximum of 400 epochs. The learning rate is initialized with a value of 0.001 and a weight decay of 5e-4. An early stopping criterion has a patience value of 20 epochs. The entire dataset is divided into 85% Train set, 10 % Test set, and 5 % Validation set and this has been maintained across all the datasets.
@@ 5 Regression
In this section, we have demonstrated our experiments on the regression. The main metrics used for quantifying the results have been presented in Table 5. A detailed description of the data-preprocessing steps of all formats of the dataset has been explained below. For a visual representation of our results parity plots are presented
@@ 5.1 Dataset Description
There are several properties such as boiling point, melting point, solubility, etc which are not readily available for use in the case of organic as well as inorganic compounds. These properties take time and resources to be found and extraction of group-contribution of these molecules is not easy to find. For novel compounds, these values are not readily available. Thus, the efficacy of the GNN model to directly extract the property from its SMILES [40] representation is accurately demonstrated based on some of the example properties.
For regression problems, we have taken the datasets used by Asheri et al [4] which contains over 24,259 molecular compounds along with their distinct properties. In order for the model. Among them we have chosen properties such as Boiling Point [4], Acentric factor [4], LogWs [4], LogP [4], Melting Point [4], LD50[4]. Apart from that we have also show-cased results on standard benchmark datasets such as the LipO (lipophilicity) dataset [52].
@@ 5.2 Regression Results
Our proposed GNN method demonstrates substantial advancements over prior benchmarks in the field. Building upon the foundational work of Asheri et al. [4] and Rittig et al [43], where R2 score of 0.85 has been reported, our revised GNN model achieved a notable enhancement in performance in the Boiling point dataset[4] both in training and testing phases. As a benchmark, in the test phase, our model showcased a significant improvement, achieving a 10.1% increase in R2 score on average over the single GNN model [43] and 3.41% increase over their ensembled GNN model [43], thus surpassing the previous state-of-the-art results.
To verify our model performance, we have stated the results of various other datasets of Asheri et al [4], and we have obtained competitive or better metrics on those sets. It is observed that the model performs extremely well on those sets where the number of data points is comparatively larger than others. This is inherently due to the fact the GNN models require a comparatively vast amount of data to properly generalize.
Table 5: Prediction Results from Graph Regression
@@
(a) Lipophilicity
(b) Boiling Point
(c) Melting Point
(d) LogP
(e) LogWs
(f) Acentric Factor
(g) Acid Dissolution Factor
(h) LD50
Figure 11: Parity plot of various datasets of [4] on their test sets as presented
On comparing the results of our model with benchmark datasets, such as Lipophilicity [52], we can see an RMSE score of 0.1 on average on the test set, which is a massive improvement from [58] and [59] which are currently the top benchmarks in the Lipophilicity dataset [52] and reports a RMSE score of 0.587 along with pretraining on additional dataset. It is to be noted that for modeling we have not used any externally added datasets which could have helped in generalization. The results of this paper not only live up to the standards of the current benchmark models but also show the robust generalization capability of our model across all sorts of datasets.
To further highlight the distribution of points and visualize the results, Parity plots of the test set have been provided for all the datasets in Figure 11 to show the grouping of the results on the test set. The actual experimental values have been plotted in the x-axis whereas the predicted values have been plotted in the y-axis. From the plots, it can be seen that the data points lie mostly within the ±10% error line, even for large datasets like Boiling Point, LogP, and LD50 as shown in Figure 11(b), 11(e), 11(h). Since none of the data points have been excluded, the parity plots show that even for cases that are outliers the model performs well and the data points lie within the acceptable limits. However, we can see that the data points for melting points are not that closely packed. This can be reflected in the R2 score as well as the model’s failure to generalize much of the dataset. Apart from the Melting Point dataset 11(c), we can say that more than 98% of the dataset falls below the 10% error range.
Additionally, our method’s success on the LogW, LogP, and Acentric factor datasets highlights its adaptability and generalizability on different datasets. For a clear comparison across all datasets as well as with current benchmark models, we have added all forms of results such as MAE, MSE, and R2 which can be associated with the dataset. A comparison of the number of the data points for the test set
@@ 6 Classification
Graph classification plays a pivotal role in the prediction of behaviours of different molecules and has various applications in the domains of drug discovery [25], the generation of new materials [60], and identifying the intrinsic properties of different compounds [24]. Given the superior performance of the new GNN model in regression scenarios,it becomes imperative to subject it to classification tasks, thereby affirming its robustness in discerning intricate patterns across data of different tasks. We perform multi-label multi-class classification, a domain where past experiments have been sparse, proving the model’s ability to handle diverse and complex tasks.
In this study, a comparative analysis has been done between a conventional methodology involving GNN feature embeddings fed into a multi-neuron sigmoid layer and a novel approach with our GNNs as a backbone to a VAE model, as expounded in Sec. 3.3. The latter method facilitates end-to-end learning of both the GNN and VAE blocks at once. A comprehensive comparison is performed to highlight the results and advantages of the novel approach over the conventional approach.
@@ 6.1 Dataset Description
For the classification setting, we have focused on those datasets that have multiple classes or sub-tasks and each data point is associated with one of those labels. With this setting in mind, popular datasets have been chosen so that we can benchmark our models against other state-of-the-art algorithms. A detailed description of the datasets used has been presented. The Tox21[52] dataset includes a set of chemical compounds examined for toxicity across 12 different biological targets. With over 8,000 compounds, it provides valuable information for predicting the toxicity and potential adverse effects of various chemicals. Another dataset such as SIDER [52] provides a helpful resource about drug side effects. It gathers info from different sources like clinical trials and medical reports. It helps check if drugs are safe and understand side effects. There are 27 tasks in the dataset which give useful insights into possible side effects of different drugs. Furthermore, we have tried our classification models on higher-dimensional label space methods like Toxcast [52] which consists of more than 8000 compounds and 617 labels. Similarly, with PCBA dataset [52], generated by high throughput screening for biologically active molecules. Like Toxcast this consists of 128-dimension label space and over 400,000 compounds.
@@ 6.2 Dataset Processing
Regrettably, several SMILE-encoded datasets earmarked for multilabel multiclass classification exhibited anomalies such as missing labels and vacant entries. While rows with empty entries could be easily excluded, coping with partially missing labels posed a more intricate challenge. Imputing these labels was deemed suboptimal, as it risked compromising the authentic intricacies of the real-world data. Consequently, the preprocessing unfolded in two distinct phases. Firstly, entries with over 50% missing labels were expunged from the dataset. Despite a reduction in data volume, this step ensured the preservation of a certain qualitative threshold, retaining crucial representative elements. Secondly, addressing the unavoidable necessity, missing data was imputed with a constant value (we opted for 0). This step, however, introduced some data corruption and exacerbated the existing class imbalance. Notably, the SIDER [52] dataset remained pristine, devoid of such defects, and was used directly. In contrast, Tox21[52] and ToxCast [52] datasets suffered more significantly from this issue.
@@
Below shown in Figure 12 is a statistical analysis of the imbalances in the dataset with respect to the labels. As can be seen, there is a huge mismatch between the 1’s and 0’s of all the datasets. One of the major problems in these datasets is the presence of huge number of NAN/missing values. This, problem although not present with sider is predominant in the other datasets such as Tox21[52], PCBA [52], and especially ToxCast [52]. Since imputing all the values with 0 was not deemed ideal, we have also conducted experiments by giving the NAN values a separate class to give them proper representation, the results of which have been discussed. Thus, our problem formulation has been presented for both binary classes as well as multi-class multi-label classification.
6.3 Results of GNN+sigmoid
Before the GNN + VAE model, ablation studies have been performed with the base GNN model. Table 6 describes the results obtained directly from the GNN backbone with the last layer being the sigmoid layer. As can be seen from the table the results are not satisfactory enough with just a single sigmoid layer. The scores obtained however surpass those obtained from the GCN [33] backbone, which further highlights the model’s effectiveness.
Table 6: Results of GNN+sigmoid method
@@ 6.3 Results of GNN+sigmoid
As can be seen from the results, the primitive method of directly producing the labels is not able to be accurately represented by the GNN model and the accuracy is very low. Thus, for multi-class multi-label settings different options of accurately representing the labels to the embeddings we have presented the GNN+VAE model. A major advantage of this is that the VAE can be directly attached to the GNN backbone.
6.4 Results of GNN+VAE model
The predominant metric for multilabel multiclass classification is the micro-AUC-ROC score, as it is robust against class imbalance in different labels. Previous studies, specifically [35] and [24], utilized a single-model-per-label approach in their work on the Tox21[52] and Sider [52] datasets. While effective for molecular property prediction and drug discovery, this approach proves impractical in high-dimensional label-space scenarios, demanding substantial time and resources. In contrast, our GNN+VAE model for multilabel multiclass classification predicts all labels simultaneously, significantly reducing the resources required for both training and inference. Notably, outperforming the state-of-the-art results reported by [24] on the sider dataset, our model achieves an AUC-ROC score of 93.26%, surpassing their reported 91% average. There is also a discrepancy to be noted. The Bio-Hect [24] model reports metrics on a subset of labels, while our model provides predictions for all labels in the dataset. On the tox21 dataset, our model achieves an AUC-ROC score of 84%, somewhat lower than [35]’s mean score of 96%, attributed to the smaller number of labels in the Tox21 dataset contributing to better learning of single-model-per-label systems. Despite this, our approach proves to be a competitive and resource-efficient alternative.
On distinct 85-5-10 train-validation-test splits, our model demonstrates robust performance on sider, achieving a 76.7% hamming accuracy. Notably, the model excels on Tox21 with an impressive 60% accuracy and a consistent 92% hamming accuracy. The observed discrepancy is due to the lower label dimension in Tox21, contributing to the model’s optimal performance in this setting.
On the PCBA dataset, our model underperforms with an AUC score of 75%, quite lower than the benchmark set by [61] at 87%. In the expansive ToxCast dataset featuring an impressive 617 task labels, our model showcases competitive performance, achieving a notable 78.2% AUC-ROC score. This result narrowly surpasses the prior benchmark established by [62]. A similar approach, the Graph Multiset pooling model for graph classification [28], which, like ours, predicts all labels simultaneously with a single model, reports a considerably lower ROC-AUC score of 65.44%.
The versatility and adaptability of our model are evident in its consistent performance across these diverse datasets. The substantial enhancements in Micro-AUC-ROC scores, particularly in Sider, and the model’s ability to provide competitive results even for a high dimensional label space like that of ToxCast, underscore not only the efficacy of our approach but also its potential for accurate and comprehensive multilabel multiclass classification tasks in various domains. This in turn affirms the robustness and generalizability of the new GNN+VAE model. While our model excels at some tasks, some other datasets have better methods to deal with them individually. The upside to this is that our model performs consistently and competitively on all SMILE-encoded classification datasets with SOTA or near-SOTA performance.
Table 7: Predictions for Multi-label Binary-class Classification

### v2

@@ 3.4 End to End Learning
Since there are two problems to be solved, regression and classification, both will have their own independent learning schemes. The entire GNN backbone can be represented as $F_{n}:G(n)\rightarrow m_{f}$, where $F_{n}$ is the GNN function that generates the molecular fingerprint/embeddings $(m_{f})$. The MLP as defined in Sec.3.2 is automatically integrated in the GNN function which allows for backpropagation. The embeddings are then used for the general downstream tasks. For the regression task, the task is defined in the format $MLP(m_{f})\rightarrow\text{output prediction}$.
@@
The multilayer perceptron is applied to molecular embeddings generated by the GNN. Since the entire updation takes place in a single setting and is bound by a single computation graph, proper end-to-end learning takes place. For the multi-class multi-label classification setting, a diagrammatic representation of the workflow has been represented in Figure 10. As can be seen, the SMILES data is first converted to a graph database by PyTorch-geometric 10, with the features of the nodes and edges being extracted by Rdkit 54. The dataset is then passed through the GNN as represented above to generate $m_{f}$.The VAE part of the execution can be represented as $VAE(m_{f})\rightarrow\hat{p}$ and $\sigma(\hat{p})\rightarrow\text{predicted vector}$.
In the current scenario, an MLP is used as the encoder and the decoder of the VAE; however, it can be easily replaced with much more sophisticated VAE methods such as 55. The VAE generates a probability vector, which will give us the predicted vector when passed through the sigmoid function. In both the regression and classification scenarios, all the entities are attached to an individual computation graph. Thus, this allows for an end-to-end manner of training where the generated molecular fingerprint is dependent on all the entities, and backpropagation is allowed throughout the computation graph.
@@ 4 Model Setup and Training
For the model setup, the ECRGNN backbone described in Sec. 3.1 is used for both the regression and classification tasks. All the hyper-parameters of the graph convolutions and the fully connected layers remain the same. In the case of regression property prediction, Huber loss, a combination of MSE and MAE, was used to account for the presence of outliers. The resultant loss generated is converted into an exponential loss function.
The cross-entropy loss function has been used in the case of Multi-class/Multi-label Classification datasets. The ECRGNN backbone, along with a sigmoid layer of one-hot encoded classes, has been used. To prevent the class imbalance from affecting the results, a self-adaptive weighted loss function controlled by Adam optimizer 56 is implemented. This prevents some classes from becoming over-dominant, thus increasing the F1 score.
The ECRGNN backbone, in conjunction with VAE, has been used to establish an end-to-end pipeline for training. The hyperparameters of the GNN backbone remain the same, and those of the VAE have been described in Sec. 3.3. To explain the ECRGNN-VAE model’s efficacy, we also present comparative results of the ECRGNN backbone with the last layer as a sigmoid activation function.
Adam optimizer and Cosine annealing scheduler with warm restarts have been used in both regression and classification cases. The batch size is 64, which is found to be the optimum, and we have run the model for a maximum of 400 epochs. The learning rate is initialized with a value of 0.001 and a weight decay of 5e-4. An early stopping criterion has a patience value of 20 epochs. The entire dataset is divided into 85% Train set, 10 % Test set, and 5 % Validation set, and this has been maintained across all the datasets.
@@ 5 Regression
This section demonstrates the experiments on the regression. The main metrics used for quantifying the results have been presented in Table 5. A detailed description of the data preprocessing steps for all formats of the dataset is given below. For a visual representation of the results, parity plots are presented
@@ 5.1 Dataset Description
There are several properties, such as boiling point, melting point, solubility, etc, which are not readily available for use in the case of organic as well as inorganic compounds. These properties take time and resources to be found, and the extraction of group contributions to these molecules is not easy. For novel compounds, these values are not readily available. Thus, the efficacy of the GNN model to directly extract the property from its SMILES 39 representation is accurately demonstrated based on some of the example properties.
For regression problems, datasets are referred from Asheri et al. 4, which contain over 24,259 molecular compounds and their distinct properties. In order for the model. Among them properties such as Boiling Point 4, Acentric factor 4, LogWs 4, LogP 4, Melting Point 4, LD504 have been chosen. Apart from that, results on standard benchmark datasets such as the LipO (lipophilicity) dataset 51 have been show-cased.
@@ 5.2 Regression Results
The ECRGNN method demonstrates substantial advancements over prior benchmarks in the field. Building upon the foundational work of Asheri et al. 4 and Rittig et al. 42, where R2 score of 0.85 has been reported, the revised GNN model has achieved a notable enhancement in performance in the Boiling point dataset4 both in training and testing phases. As a benchmark, in the test phase, the ECRGNN showcased a significant improvement, achieving a 10.1% increase in R2 score on average over the single GNN model 42 and 3.41% increase over their ensembled GNN model 42, thus surpassing the previous state-of-the-art results.
Table 5: Prediction Results from Graph Regression.
@@
To verify the ECRGNN’s performance, results of various other datasets of Asheri et al. 4 are cross-referenced, and the ECRGNN obtained competitive or better metrics on these sets. It is observed that the model performs extremely well on those sets where the number of data points is comparatively larger than others. This is inherently due to the fact that GNN models require a comparatively vast amount of data to properly generalize.
Comparing the results of the ECRGNN on benchmark datasets, such as Lipophilicity 51, an RMSE score of 0.1 is obtained on average on the test set, which is a massive improvement from 57 and 58 which are currently the top benchmarks in the Lipophilicity dataset 51 and reports a RMSE score of 0.587 along with pretraining on additional dataset. It is to be noted that no externally added supplementary data has been used for modelling, which could have helped in generalization. The results of this paper not only live up to the standards of the current benchmark models but also show the robust generalization capability of the model across all sorts of datasets.
Figure 11: Parity plot of various datasets of 4 on their test sets as presented - (a) Lipophilicity (b) Boiling Point (c) Melting Point (d) LogP (e) LogW (f) Accentric Factor (g) Acid Dissolution factor (h) LD50 dataset.
To further highlight the distribution of points and visualize the results, Parity plots of the test set have been provided for all the datasets in Figure 11 to show the grouping of the results on the test set. The actual experimental values have been plotted in the x-axis whereas the predicted values have been plotted in the y-axis. From the plots, it can be seen that the data points lie mostly within the ±10% error line, even for large datasets like Boiling Point, LogP, and LD50, as shown in Figure 11(b), 11(e), 11(h). Since none of the data points have been excluded, the parity plots show that even for cases that are outliers, the model performs well, and the data points lie within the acceptable limits. However, it can be seen that the data points for melting points are not that closely packed. This can be reflected in the R2 score as well as the model’s failure to generalize much of the dataset. Apart from the Melting Point dataset 11(c), more than 98% of the dataset for other datasets fall below the 10% error range.
Additionally, the method’s success on the LogW, LogP, and Acentric factor datasets highlights its adaptability and generalizability on different datasets. For a clear comparison across all datasets as well as with current benchmark models, all forms of results such as MAE, MSE, and R2 scores have been recorded for each dataset.
@@ 6 Classification
Graph classification plays a pivotal role in the prediction of behaviours of different molecules and has various applications in the domains of drug discovery 35, the generation of new materials 59, and identifying the intrinsic properties of different compounds 33. Given the superior performance of the new GNN model in regression scenarios, subjecting it to classification tasks is imperative, thereby affirming its robustness in discerning intricate patterns across data of different tasks. Multi-label multi-class classifications are performed. It is a domain where past experiments have been sparse, proving the model’s ability to handle diverse and complex tasks.
In this study, a comparative analysis has been done between a conventional methodology involving GNN feature embeddings fed into a multi-neuron sigmoid layer and a novel approach with the GNNs as a backbone to a VAE model, as expounded in Sec. 3.3. The latter method facilitates end-to-end learning of the GNN and VAE blocks simultaneously. A comprehensive comparison is performed to highlight the results and advantages of the novel approach over the conventional approach.
@@ 6.1 Dataset Description
For the classification setting, the focus has been on datasets with multiple classes or sub-tasks, and each data point is associated with one of those labels. With this setting in mind, popular datasets have been chosen to benchmark the ECRGNN’s performance against other state-of-the-art algorithms. A detailed description of the datasets used has been presented. The Tox2151 dataset includes a set of chemical compounds examined for toxicity across 12 different biological targets. With over 8,000 compounds, it provides valuable information for predicting the toxicity and potential adverse effects of various chemicals. Another dataset such as SIDER 51 provides a helpful resource about drug side effects. It gathers info from different sources like clinical trials and medical reports. It helps check if drugs are safe and understand side effects. There are 27 tasks in the dataset that provide useful insights into the possible side effects of different drugs. Furthermore, the classification models have been tried on higher-dimensional label space methods like Toxcast 51, which consists of more than 8000 compounds and 617 labels. Similarly, with PCBA dataset 51, generated by high throughput screening for biologically active molecules. Like Toxcast, this has a 128-dimensional label space and over 400,000 compounds.
@@ 6.2 Dataset Processing
Regrettably, several SMILE-encoded datasets earmarked for multilabel multiclass classification exhibited anomalies such as missing labels and vacant entries. While rows with empty entries could be easily excluded, coping with partially missing labels posed a more intricate challenge. Imputing these labels was deemed suboptimal, as it risked compromising the authentic intricacies of the real-world data. Consequently, the preprocessing unfolded in two distinct phases. Firstly, entries with over 50% missing labels were expunged from the dataset. Despite a reduction in data volume, this step ensured the preservation of a certain qualitative threshold, retaining crucial representative elements. Secondly, missing data was imputed with a constant value (initially 0) to address the unavoidable necessity. This step, however, introduced some data corruption and exacerbated the existing class imbalance. Notably, the SIDER 51 dataset remained pristine, devoid of such defects, and was used directly. In contrast, Tox2151 and ToxCast 51 datasets suffered more significantly from this issue.
@@
Figure 12 is a statistical analysis of the imbalances in the dataset with respect to the labels. As can be seen, there is a huge mismatch between the 1’s and 0’s of all the datasets. One of the major problems in these datasets is the presence of a huge number of NAN/missing values. This problem, although not present with sider, is predominant in the other datasets such as Tox2151, PCBA 51, and especially ToxCast 51. Since imputing all the values with 0 was not deemed ideal, experiments were conducted by giving NAN values a separate class to give them proper representation, the results of which have been discussed. Thus, the problem formulation has been presented for both binary classes as well as multi-class multi-label classification.
6.3 Results of ECRGNN+sigmoid
Before the ECRGNN + VAE model, ablation studies were performed using the base ECRGNN model. Table 6 describes the results obtained directly from the ECRGNN backbone, with the last layer being the sigmoid layer. As can be seen from the table, the results are not satisfactory enough with just a single sigmoid layer. The scores obtained, however, surpass those obtained from the ECRGCN 30 backbone, which further highlights the model’s effectiveness.
Table 6: Results of GNN+sigmoid method.
@@ 6.3 Results of ECRGNN+sigmoid
As can be seen from the results, the primitive method of directly producing the labels is not able to be accurately represented by the ECRGNN model, and the accuracy is very low. Thus, for multi-class multi-label settings, different options for accurately representing the labels to the embeddings, the ECRGNN+VAE model, has been introduced. A major advantage of this is that the VAE can be directly attached to the ECRGNN backbone.
6.4 Results of ECRGNN+VAE model
The predominant metric for multilabel multiclass classification is the micro-AUC-ROC score, as it is robust against class imbalance in different labels. Previous studies, specifically Elembert-v132 and BioAct-Het33, utilized a single-model-per-label approach in their work on the Tox2151 and Sider 51 datasets. While effective for molecular property prediction and drug discovery, this approach proves impractical in high-dimensional label-space scenarios, demanding substantial time and resources. In contrast, the GNN+VAE model for multilabel multiclass classification predicts all labels simultaneously, significantly reducing the resources required for both training and inference. Notably, outperforming the state-of-the-art results reported in Table 7 on the sider dataset, the model achieves an AUC-ROC score of 93.26%, surpassing their reported 91% average. There is also a discrepancy to be noted. The BioAct-Het 33 model reports metrics on a subset of labels, while ECRGNN provides predictions for all labels in the dataset. On the tox21 dataset, ECRGNN achieves an AUC-ROC score of 84%, somewhat lower than Elembert’s32 mean score of 96%, attributed to the smaller number of labels in the Tox21 dataset contributing to better learning of single-model-per-label systems. Despite this, the approach proves to be a competitive and resource-efficient alternative.
Table 7: Predictions for (Imputation 0) Classification.